Suppose that you invest $7,000 today at an annual rate of 9%. Assuming that the expected annual rate of inflation is 4%, what will be your increase in purchasing power in percentage terms? Enter your answer as a percent without the “%.” Round your final answer to two decimals.
Answer #1

Nominal increase in our investment = 7,000 * 0.09 = $630

Inflation takes a part of it = 7,000 * 0.04 = $280

So, our increase in purchasing power = 630 - 280 = $350 (roughly)

In percentage terms, our increase in purchasing power 350/7000 = 0.05 = 5% (roughly)

Actual calculations:

(1 + real rate) = (1 + nominal rate)/(1 + inflation)

(1 + real rate) = (1 + 0.09)/(1 + 0.04)

(1 + real rate) = 1.0480769231

real rate = 1.0480769231 - 1

real rate = 0.0480769231

real rate = 4.80769231%

Our purchasing power increases by 4.80769231%.

Can you please upvote? Thank You :-)